# Enhancing Robot Safety via MLLM-Based Semantic Interpretation of Failure Data

### v1

@@ 5.1 Case Study 1: Real-World Car Crash Analysis from Dashcam Videos
We apply our framework to the Nexar driving dataset [30], which includes 1,500 dashcam videos (each approximately 40 seconds, 1280×720 resolution at 30 fps) of collisions or near-misses involving the ego vehicle. While these recordings are from human-driven vehicles, they serve as a proxy for autonomous driving failures in the absence of large-scale public AV failure datasets. However, our framework is directly applicable to autonomous vehicle logs as well.
5.1.1 Failure Cluster Discovery
@@
Our system successfully discovers a diverse set of interpretable failure clusters from the driving dataset, as shown in Fig. 3. Qualitative inspection (Fig. 4) shows that these clusters correspond to meaningful and recurring traffic incident types, such as rear-end collisions, unsafe lane changes or intersection misjudgments.
@@
Baselines and Results.
To evaluate the quality of the failure clusters discovered by our method, we compare against BERTopic [21], a state-of-the-art topic modeling method that combines transformer embeddings with unsupervised clustering and keyword extraction. We apply BERTopic to the same set of failure reasons produced in Step 1 of our pipeline to ensure a direct and fair comparison. We also evaluate a stronger variant, BERTopic-LLM, where a language model summarizes each discovered topic using representative keywords and samples.
Figures 5 and 6 illustrate the clusters and keywords produced by BERTopic and BERTopic-LLM, respectively. Standard BERTopic struggles to generate semantically coherent and interpretable clusters. Key failure modes — e.g., Rear-End Collisions— are diluted across vague or syntactically noisy topics (e.g., lead_vehicle_to_failed, white_the_lane_ego). BERTopic-LLM partially improves interpretability, recovering some valid modes such as Rear-end collision mechanics, but still includes vague or tautological categories such as Driving safety violations, which do not add diagnostic value or clearly differentiate failure modes. In contrast, our method consistently yields sharper, failure-relevant clusters (Fig. 3).
5.1.2 Failure Monitoring Leveraging Discovered Clusters
We evaluate runtime failure detection by prompting a multimodal reasoning model with recent image history and the learned failure clusters and reasoning about any possible near-future collision. Refer Appendix A.1.2 for more details.
Baselines.
We compare our method against VLM-Based Anomaly Detection (VLM-AD) methods [17]. These methods provide scene descriptions to an LLM and ask it to detect any possible anomalies in the current observation. We also compare against Leaderboard Method – the top-performing failure video classifier on the Kaggle Nexar challenge, based on fine-tuned VideoMAEv2-giant [42] Refer Appendix A.1.2 for details..
Results.
We compare all methods on a set of held-out driving trajectories. As shown in Table 1, our method consistently achieves the highest F1 score and outperforms baselines in both True Positive Rate (TPR) and False Negative Rate (FNR), showcasing its ability to detect the actual failures both robustly and reliably. These results also indicate that the system failures are not always the same as anomalies or out-of-distribution inputs; in-distribution scenarios, like those specified in our failure clusters, can also lead to system failures that are hard to capture with VLM-AD. We also test all methods on an unseen dashcam dataset of 200 trajectories (see Table 1 (Middle)). The proposed method demonstrates strong generalization. This suggests our method captures structured semantic patterns beyond dataset-specific cues, unlike the leaderboard classifier, which lacks generalizability.
Finally, we compare the lead failure detection time of different methods (Table 1 (Right)). Our method detects failures earlier than others, indicating a stronger ability to anticipate failures by correlating the scene observation with the failure clusters.
In-Distribution Trajectories | Out-of-Distribution Trajectories | ||||||||||
Method | TPR | TNR | FPR | FNR | F1 | TPR | TNR | FPR | FNR | F1 | Time |
Ours | 71.4 | 72.5 | 27.5 | 28.6 | 71.4 | 83.0 | 70.0 | 30.0 | 17.0 | 77.9 | 610 |
NoContext | 42.8 | 85.3 | 14.7 | 57.2 | 54.1 | 64.0 | 80.0 | 20.0 | 36.0 | 69.6 | 473.3 |
VLM-AD | 7.1 | 91.1 | 8.9 | 92.9 | 12.3 | 35.0 | 94.0 | 6.0 | 65.0 | 49.7 | 166.6 |
Leaderboard | 52.0 | 93.1 | 6.9 | 48.0 | 65.3 | 18.0 | 75.0 | 25.0 | 82.0 | 25.2 | 506.6 |
Ablations.
We also compare the performance of the proposed failure detection method on removing the failure cluster information from the prompt (NoContext). All other implementation details, such as processing a history of past observations and CoT reasoning, remain the same as our method. This results in significant performance degradation, highlighting the utility of failure clusters information in better detection across environments.
5.2 Case Study 2: Failure Discovery in Indoor Robot Navigation
@@
We apply our framework to a vision-based ground robot, navigating unknown indoor office environments [4]. The robot uses a CNN-based policy with a model-based low-level controller. It receives RGB images, ego-velocities, and a goal position, and outputs acceleration commands for the robot. We record robot rollouts in the Stanford office environment [3] and extract front-view image sequences. Colliding trajectories comprise our failure dataset , which we use for clustering.
5.2.1 Failure Cluster Discovery
@@
Our method discovers a set of interpretable failure clusters from the collision trajectories (Fig. 7). Notably, clusters such as walls and chairs were previously identified by manual inspection and reachability analysis in [10]. This validates our method’s ability to automatically recover known failure types and uncover new semantic patterns.
5.2.2 Targeted Data Collection and Policy Fine-Tuning for Enhanced Safety
@@
Following Sec. 4.2, we use the discovered clusters to guide expert data collection in targeted regions of the environment. The robot policy is fine-tuned on this augmented dataset (details in Appendix A.2.3). The failure rate in sampled trajectories drops from 46% to 18%, demonstrating enhanced safety in previously failure-prone situations. This forms a closed-loop pipeline of failure discovery, targeted intervention, and policy refinement for continuously enhancing system safety.
5.2.3 Failure Monitoring
@@
We use the same runtime monitoring approach as in Case Study 1. The monitor reasons over the scene, past trajectory, and known failure clusters to preemptively identify any potential failure. As evident in Fig. 8, due to the confined and cluttered nature of the indoor environment, the failure monitor can easily misinterpret static background elements with prior failure contexts, leading to a high false positive rate. To mitigate this, we introduce a simple temporal consistency rule: a failure is flagged only if it persists for three consecutive frames. This helps reduce conservativeness while preserving responsiveness. Refer Appendix A.2.2 for details.
Baselines.
@@
We compare against VLM-AD, where we adapt the prompt from [17] by giving examples relevant to an indoor office environment in the prompt. We additionally compare against ENet-BC, a vision-based runtime failure monitor based on EfficientNet-B0, trained on labeled collision data from the same environment [10]. Note that, unlike our method and VLM-AD, ENet-BC requires environment-specific training and does not generalize to unseen layouts (details in Appendix A.2.2).
Results.
On an in-distribution (IID) test set of 326 trajectories, our method outperforms all LLM-based baselines in F1 score (Table 2, Left). ENet-BC achieves similar performance to the proposed method, as expected given its environment-specific training.
In-Distribution Trajectories | Out-of-Distribution Trajectories | ||||||||||
Method | TPR | TNR | FPR | FNR | F1 | TPR | TNR | FPR | FNR | F1 | Time |
Ours | 65.0 | 99.0 | 1.0 | 35.0 | 77.2 | 67.6 | 86.2 | 13.8 | 32.4 | 50.0 | 1.21 |
NoContext | 51.7 | 99.6 | 0.4 | 48.3 | 67.37 | 45.9 | 89.0 | 10.1 | 54.1 | 40.5 | 0.76 |
VLM-AD | 83.3 | 50.0 | 50.0 | 16.7 | 40.0 | 62.2 | 60.1 | 39.9 | 37.8 | 27.2 | 1.38 |
ENet-BC | 65.0 | 100.0 | 0.0 | 35.0 | 78.8 | 100.0 | 6.3 | 93.7 | 0.0 | 22.4 | 1.01 |
@@ Results.
To test generalization, we evaluate on an out-of-distribution (OOD) set of 300 trajectories from a different building. The performance of all methods degrades as expected, but the proposed method maintains the highest F1 score, while ENet-BC fails entirely to generalize. This again demonstrates the generalization capabilities of our method.
@@
Our method also detects failures earlier on average, highlighting its ability to reason about impending collisions before impact. Full metrics are in Table 2. Although VLM-AD has a slightly higher average time, it is because it fails entirely and detects everything as anomalous, evident from its high FPR.
@@
We integrate our failure monitor with a reactive safeguard controller that activates upon a failure detection. Fig. 8 shows an example where the nominal policy leads to a collision, while the monitor detects the failure and invokes the safeguard controller, enabling successful recovery.
6 Conclusion and Limitations
@@
We present a closed-loop framework for interpretable failure analysis in autonomous systems. Our method automatically discovers semantically meaningful failure modes from perception recordings without supervision, distilling them into human-understandable clusters annotated with natural language. These clusters support targeted data collection, policy fine-tuning, and semantic failure detection—enabling scalable and continuous safety improvement. By leveraging the reasoning capabilities of MLLMs, our approach provides a foundation for understanding, organizing, and responding to failures in complex, unstructured environments.
@@
While our framework enables interpretable and actionable failure analysis, it has several limitations. First, there is no canonical “correct” way to cluster failure trajectories. Different clustering strategies may highlight different operational insights. Future work could integrate formal techniques like STPA (Systems-Theoretic Process Analysis) or FRAM (Functional Resonance Analysis Method) to complement unsupervised clustering. Second, MLLMs can generate plausible-sounding but incorrect failure explanations, or miss the true cause entirely. Future work could incorporate simulation-based counterfactual testing or causal world models to verify and refine inferred explanations. Finally, while we demonstrate results on datasets of 700 trajectories, our pipeline can scale to larger datasets and temporally extended events with appropriate batching and model throughput optimization.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge research support from DARPA Assured Neuro Symbolic Learning and Reasoning (ANSR) program and by the NSF CAREER program (2240163).
Appendix A Technical Appendices and Supplementary Material
The appendix offers additional details, such as prompts, about our implementation. Sections A.1 and A.2 provide the details for dashcam driving datasets and indoor ground navigation, respectively.
A.1 Case Study 1: Real-World Car Crash Analysis from Dashcam Videos
Section A.1.1 contains all the details of discovering the failure clusters, and Section A.1.2 lists all the details about runtime monitoring for the driving case study. The grey boxes contain prompts.
A.1.1 Failure Cluster Discovery
Step 1: Inferring Failure Reasons with MLLMs.
Step 2: Discovering Semantic Failure Clusters.
We generate an ensemble of prompts from an LLM (OpenAI o4-mini) to infer multiple failure clustering results, which are then aggregated into the final list of clusters. Here are the three prompts generated by the LLM:
Prompt 1:
Prompt 2:
Prompt 3:
Each of these prompts is appended with a list of all the failure reasons from the previous step. Finally, we query the LLM again to consolidate these individual lists and generate the final clusters.
Step 3: Assigning Trajectories to Failure Clusters.
We compare the distribution of discovered clusters to their actual assignments across trajectories (Fig. 9). Most frequencies align closely. The largest discrepancies appear for the visibility and sudden intrusion clusters. Visibility failures show higher reported than assigned frequency in single-label settings, but match under multi-label assignment—suggesting co-occurrence with other failures. Similarly, sudden intrusions is under-assigned due to ambiguity with intersection or cut-in cases.
A.1.2 Failure Monitoring Leveraging Discovered Clusters
As described in Sec. 4.2, we directly use the obtained failure clusters to safeguard the system through runtime monitoring. The reasoning model takes a history of past image observations as input and tries to reason about any possible near-future collision while leveraging the knowledge about the most common failures of that system from given clusters in a Chain-of-Thought (CoT) manner. Adding the cluster information in the runtime monitor’s prompt helps in zero-shot failure detection for timely executing a safe action, and generalizing across different agents and environments without requiring any specialized training. Here is the prompt:
The prompt first outlines the main task for the LLM to act as a runtime failure monitor and detect if the system is currently SAFE or UNSAFE based on the observation history. Further, it provides more information about the system, such as what is meant by a failure here and the frequency of observations. Then, we ask it to sequentially reason about the possible future motion of the system, the surrounding environment, and any possible risk of failure given the most common failure modes of that system through clusters. list of clusters contains all the generated clusters along with their keywords and descriptions.
We report the failure detection metrics on both In-Distribution (IID) Trajectories and Out-of-Distribution Trajectories (OOD) in Table 1. For IID testing, we took a held-out set of 200 trajectories from the same Nexar dataset, which we used for clustering analysis. For OOD testing, we took another open-source dashcam dataset [5] containing ego car crash videos.
Fig. 10 shows a few examples for True Positives and False Negatives predictions from our method. Rows 1,2 are successfully detected as unsafe. Rows 3,4 are failures, but are detected as safe because the failure event is not evident in those images. Note that the last image in the sequence is taken 1 sec before the collision, and since car driving is a high-speed system, the actual failure sometimes starts even after that, which leads to false safe predictions.
Baselines.
VLM-AD is an LLM-based few-shot runtime anomaly detection method, where the current image observation is first processed through a scene descriptor, such as an object detector, which further feeds into an LLM. The LLM is asked to find any possible anomalous scenario in the current scene, by providing it with a description of some Normal and Anomalous examples in the prompt. For the driving case study, we directly take the prompt from [17] and use the same reasoning LLM as our method to prevent any model-specific performance difference and perform a fair comparison of the methods. It should also be noted that these methods require manually writing few-shot examples in the prompt to provide contextual information about the system, which in our method is automated with the generated clusters.
For the Leaderboard-Method baseline, since the Nexar dataset is also hosted as a Crash Prediction Challenge on Kaggle, we take the best publicly available method [1] as per leaderboard scores to compare with our method. It takes a pre-trained VideoMAEv2-giant model as the backbone and trains a Linear binary classification MLP layer to predict crash. VideoMAEv2-giant [42] is trained on a large amount of video datasets, such as action recognition datasets, sports videos, instructional videos, etc. We tested their trained model on our test dataset for both the IID and OOD cases. Being a classical vision-based classification approach, this method requires environment-specific training and fails to generalize.
A.2 Case Study 2: Failure Discovery in Indoor Robot Navigation
Section A.2.1 contains all the details of discovering the failure clusters, and Section A.2.2 and Section A.2.3 contain all the details about runtime monitoring and retraining for the indoor robot navigation, respectively. The grey boxes contain prompts.
A.2.1 Failure Cluster Discovery
Step 1: Inferring Failure Reasons with MLLMs.
We randomly sample initial states in the environment and trajectory rollouts with a fixed goal location. For each rollout, we record the front-view image observations and a binary label denoting Failure or Success. Further, we filter out the failure trajectories for our purpose. The following clustering analysis is performed on 228 failure trajectories.
Step 2: Discovering Semantic Failure Clusters.
Here is the ensemble of prompts given by the LLM. Each of these prompts is appended with a list of all the failure reasons from the previous step.
Prompt 1:
Prompt 2:
Prompt 2:
Prompt 3:
Step 3: Assigning Trajectories to Failure Clusters.
We compare the distribution of discovered clusters to their actual assignments across trajectories (Fig. 9). Most frequencies align closely. The largest discrepancies appear for the Narrow-Gap/Clearance Misjudgment and Low-Height Clutter & Small Floor Obstacles clusters.
They show higher reported than assigned frequency in single-label settings, but match under multi-label assignment–suggesting co-occurrence with other failures. Similarly, Bins & Waste Receptacles is under-assigned due to ambiguity with Low–Height Clutter & Small Floor Obstacles or Box--Like Equipment & Carts.
A.2.2 Failure Monitoring Leveraging Discovered Clusters
We perform runtime failure detection in the same way as described in A.1.2 for the indoor navigation robot. The prompt is tailored to an indoor robot with its list of clusters. Here is the detailed prompt:
We report the failure detection metrics on both IID and OOD test sets 2. The IID test set was taken from the same environment, but the OOD test set was taken from a different building with very different semantics, but an indoor office environment in the dataset.
Fig. 12 shows a few examples for True Positives and False Negatives predictions from our method. Rows 1,2 are successfully detected as unsafe. Rows 3,4 are failures, but are detected as safe because the monitor is getting confused seeing some free space ahead, and the temporal consistency rule gives false safe outputs.
Baselines.
For the VLM-AD baseline, we adapt the prompt from [17] for an indoor robot and test it in the same way as described in A.1.2. Here is the prompt:
For the ENet-BC baseline, we follow [9] and first compute the Backward Reach-Avoid Tube (BRAT) for our environment. Further, we collect a dataset of 3000 images with 1500 safe and 1500 unsafe images labeled using the BRAT and follow [22] to train the EfficientNet-B0 model for binary classification. Similar to the Leaderboard method in the driving example, it is based on environment-specific training, and there’s no opportunity for generalization across environments, where LLM-based methods shine.
A.2.3 Targeted Data Collection and Policy Fine-Tuning for Enhanced Safety
For targeted data collection around the mined failure modes of the robot, we identify regions in the environment corresponding to each cluster. In each failure region, for some starting states and a goal state, we use a Model Predictive Control (MPC) scheme to find a sequence of dynamically feasible waypoints [4]. The collected dataset around all the failure modes comprised approximately 40K training samples, which were then augmented with the original training data. Lastly, we fine-tuned the trained checkpoint for only 20 iterations on the augmented dataset and compared the performance on the previous policy and the retrained policy by rolling out trajectories for 50 initial states in the same environment.
Fig. 13 shows a few examples from the targeted dataset.
Generated clusters with keywords and their descriptions for driving dashcam datasets.
Generated clusters with keywords and their descriptions for indoor robot navigation.

### v2

@@ V-A2 Case Study 2: Real-World Car Crash Videos
We apply our framework to the Nexar car crash dataset [38], which includes 1,500 dashcam videos (each approximately 40 seconds, 1280×720 resolution at 30 fps) of collisions or near-misses involving the ego vehicle. While these recordings are from human-driven vehicles, they serve as a proxy for autonomous driving failures in the absence of large-scale public AV failure datasets. However, our framework is directly applicable to autonomous vehicle logs as well.
@@
Our system successfully discovers a diverse set of interpretable failure clusters from the driving dataset, as listed above and in Fig. 6. Qualitative inspection shows that these clusters correspond to meaningful and recurring traffic incident types, such as rear-end collisions, unsafe lane changes or intersection misjudgments.
@@
V-A3 Case Study 3: Vision-Based Indoor Robot Navigation
@@
We apply our framework to a vision-based ground robot, navigating unknown indoor office environments [40]. The robot uses a CNN-based policy with a model-based low-level controller. It receives RGB images, ego-velocities, and a goal position, and outputs acceleration commands for the robot. We record robot rollouts in the simulated Stanford office environment [41] and extract front-view image sequences. The trajectories resulting in a collision comprise our failure dataset , which we use for clustering.
@@
Our method discovers a set of interpretable failure clusters from collision trajectories, as listed below and in Fig. 7. Notably, clusters such as protruding corners, white walls, and glass doors were also previously identified by HJ-reachability analysis and manual inspection in [42]. This validates our method’s ability to automatically recover known failure types and uncover new semantic patterns.
V-B Downstream Safety Enhancement
We present the results for runtime failure monitoring for the car crash videos and the vision-based indoor navigation systems. We also present targeted data collection and policy fine-tuning results on the latter, as we have access to its operating policy, while we only have a dataset for the former.
V-B1 Case Study 1: Real-World Car Crash Videos
| Method | In-Distribution Trajectories | Out-of-Distribution Trajectories | Time | ||||||||
| TPR | TNR | FPR | FNR | F1 | TPR | TNR | FPR | FNR | F1 | ||
| Real-World Car Crash Videos | |||||||||||
| VideoMAE-BC [43] | 52.0 | 93.1 | 6.9 | 48.0 | 65.3 | 18.0 | 75.0 | 25.0 | 82.0 | 25.2 | 506.6 ms |
| LLM-AD | 7.1 | 91.1 | 8.9 | 92.9 | 12.3 | 35.0 | 94.0 | 6.0 | 65.0 | 49.7 | 166.6 ms |
| NoContext | 42.8 | 85.3 | 14.7 | 57.2 | 54.1 | 64.0 | 80.0 | 20.0 | 36.0 | 69.6 | 473.3 ms |
| Ours | 71.4 | 72.5 | 27.5 | 28.6 | 71.4 | 83.0 | 70.0 | 30.0 | 17.0 | 77.9 | 610 ms |
| Vision-Based Indoor Robot Navigation | |||||||||||
| ENet-BC [34] | 65.0 | 100.0 | 0.0 | 35.0 | 78.8 | 100.0 | 6.3 | 93.7 | 0.0 | 22.4 | 1.01 sec |
| LLM-AD | 83.3 | 50.0 | 50.0 | 16.7 | 40.0 | 62.2 | 60.1 | 39.9 | 37.8 | 27.2 | 1.38 sec |
| NoContext | 51.7 | 99.6 | 0.4 | 48.3 | 67.4 | 45.9 | 89.0 | 10.1 | 54.1 | 40.5 | 0.76 sec |
| Ours | 65.0 | 99.0 | 1.0 | 35.0 | 77.2 | 67.6 | 86.2 | 13.8 | 32.4 | 50.0 | 1.21 sec |
Failure Monitoring Leveraging Discovered Clusters
We evaluate online failure detection by prompting an MLLM with recent visual history and the discovered failure clusters, and reasoning about any possible near-future collision.
Baselines
We compare our approach against LLM-Based Anomaly Detection (LLM-AD) methods, such as [44], which provide a few human-written examples of anomalous and nominal objects for a particular system to an LLM, along with the scene description, and ask it to detect any possible anomalies in the current scene. We also compare against VideoMAE-BC, the top-performing failure video classifier on the Nexar Crash Prediction Challenge on Kaggle, based on a fine-tuned VideoMAEv2-giant model[43].
Results
We compare all methods on a set of held-out in-distribution (IID) driving trajectories. Table II shows that our method consistently achieves the highest F1 score and outperforms baselines in both True Positive Rate (TPR) and False Negative Rate (FNR), showcasing its ability to detect the actual failures both robustly and reliably. These results also indicate that the system failures are not always the same as anomalies or out-of-distribution (OOD) inputs; IID scenarios, like those specified in our clusters, can also lead to system failures that are hard to capture with LLM-AD. We also test all methods on an unseen dashcam dataset of 200 OOD trajectories. The proposed method demonstrates strong generalization. This suggests our method captures structured semantic patterns beyond dataset-specific cues, unlike VideoMAE-BC, which lacks generalizability.
Finally, we compare the lead failure detection time of different methods in Table II. Our method detects failures earlier than others, indicating a stronger ability to anticipate failures by correlating the scene observation with the clusters.
Ablations
We also compare the performance of the proposed failure detection method on removing the failure cluster information from the runtime monitor’s prompt (NoContext). All other implementation details, such as processing a history of past observations and CoT reasoning, remain the same as our method. This results in significant performance degradation, highlighting the utility of failure cluster information in better detection across environments.
V-B2 Case Study 2: Vision-Based Indoor Robot Navigation
Targeted Data Collection and Policy Fine-Tuning
@@ V-B2 Case Study 2: Vision-Based Indoor Robot Navigation
Following Sec. IV-B, we use the discovered clusters to guide expert data collection in targeted regions of the environment. The robot policy is fine-tuned on an augmented dataset containing an additional 40K samples collected in identified failure zones along with the original training data. The failure rate in sampled trajectories drops from 46% to 18%, demonstrating enhanced safety in previously failure-prone situations, whereas fine-tuning with randomly collected additional data only improves the failure rate to 34%. This forms a closed-loop pipeline of failure discovery, targeted intervention, and policy refinement for continuously enhancing system safety.
Failure Monitoring
@@
We use the same runtime monitoring approach as in Case Study 1. The monitor reasons over the scene, past trajectory, and known failure clusters to preemptively identify any potential failure. As evident in Fig. 7, due to the confined and cluttered nature of the indoor environment, the failure monitor can easily misinterpret static background elements with prior failure contexts, leading to a high false positive rate. To mitigate this, we introduce a simple temporal consistency rule: a failure is flagged only if it persists for three consecutive frames. This helps reduce conservativeness while preserving responsiveness.
Baselines
@@
We compare against LLM-AD by providing nominal and anomalous examples relevant to an indoor office environment in the prompt. We additionally compare against ENet-BC, a vision-based binary failure classifier based on EfficientNet-B0, trained on labeled collision data from the same environment [34]. Note that, unlike our method and LLM-AD, ENet-BC requires environment-specific training and does not generalize to unseen layouts.
Results
@@
On an IID test set of 326 trajectories, our method outperforms all LLM-based baselines in F1 score, as presented in Table II. ENet-BC achieves similar performance to the proposed method, as expected given its environment-specific training. To test generalization, we evaluate on an OOD set of 300 trajectories from a different building. The performance of all methods degrades as expected, but the proposed method maintains the highest F1 score, while ENet-BC fails entirely to generalize. This again demonstrates the generalization capabilities of our method.
@@
Our method also detects failures earlier on average, highlighting its ability to reason about impending collisions before impact. Full metrics are in Table II. Although LLM-AD has a slightly higher average time, it is because it fails entirely and detects everything as anomalous, as evidenced by its high FPR.
@@
We integrate our runtime failure monitor with a reactive safeguard controller that activates upon failure detection. Fig. 8 shows an example where the nominal policy leads to collision with a glass door, while the monitor proactively detects the failure and invokes the safeguard controller, enabling successful recovery.
VI Conclusion
@@
We present a closed-loop framework for interpretable failure analysis in autonomous systems. Our method automatically discovers semantically meaningful failure modes from perception recordings without supervision, organizing them into human-understandable clusters. These clusters enable targeted data collection, policy fine-tuning, and semantic failure detection—supporting scalable and continuous safety improvement. By leveraging the reasoning capabilities of MLLMs, our approach provides a foundation for understanding and responding to failures in unstructured environments.
@@
While our framework enables interpretable and actionable failure analysis, several limitations remain. There is no canonical way to cluster failure trajectories, and different strategies may highlight different insights. Future work could integrate formal methods like STPA or FRAM to complement unsupervised clustering. MLLMs may also produce plausible but incorrect explanations, which could be mitigated through causal or simulation-based validation. Finally, while we evaluate on datasets of 700 trajectories, our pipeline can readily scale to larger and more temporally extended data.